OpenStudy (soahelp):

Which of the following is equivalent to (ab)square4? i think its c a4b a4b4 ab4 4ab

OpenStudy (unklerhaukus):

\[\large (ab)^n=a^nb^n\]

OpenStudy (soahelp):

oh okay i see y because for (ab) square 4 it is actually a*4 B*4 a+b= whatever right

OpenStudy (unklerhaukus):

\[(ab)^4=(ab)(ab)(ab)(ab)=abababab=aaaabbbb=a^4b^4\]

OpenStudy (soahelp):

okay thankfully ik what your saying

OpenStudy (soahelp):

thxs for your help heres a medal